OpenStudy (anonymous):

Write an equation of a circle with the center (-5,-8) and radius 2. a.(x-5)^2+(y-8)^2=2 b.(x+5)^2+(y+8)^2=4 c.(x-5)^2+(y-8)^2=4 d. (x+5)^2+(y-8)^2=2

OpenStudy (hoblos):

the general formula is (x-a)^2 + (y-b)^2 = r^2 where (a,b) is center and r is radius so here (x+5)^2+(y+8)^2=4

OpenStudy (anonymous):

Oh alright thanks, I forgot about the gen formula, havent done math in a long time

OpenStudy (hoblos):

glad to help :)